# AutoLibra: Agent Metric Induction from Open-Ended Feedback

### v1

@@ 3.3 How aligned are the steps in AutoLibra with human judgment?
Grounding 0.95 0.95 0.98 0.93 0.93 0.95 (±$\pm 0.03$)
LLM-as-a-Judge 0.90 0.85 0.95 1.00 0.90 0.92 (±$\pm 0.04$)
Meta-Evaluation 0.98 0.90 0.85 0.83 0.95 0.90 (±$\pm 0.04$)
@@ 4 AutoLibra as a lens [Uncaptioned image]: agent evaluation with AutoLibra
Table 2: Comparison between AutoLibra-induced metrics and expert proposed evaluation dimensions and failure categories. The percentages in parenthesis denote frequency of -1 in AutoLibra-induced metrics, and failure frequency or score from the original papers.
CoGym
@@
In Sotopia, the authors (Zhou et al., 2024b) proposed 7 dimensions for evaluating social intelligence in AI agents. With AutoLibra, we recover the exact dimension Goal Completion, and 3 metrics as the subdimensions of Believability, indicating that Believability could be too high-level, while AutoLibra provides more concrete breakdown metrics. The failure rate (frequency of a score of -1 metric rating, indicating the agent performs poorly on that metric) measured by AutoLibra in these two categories roughly matches the score of the Sotopia dimensions of the agent we studied. AutoLibra induces another four metrics that are overlooked in the heuristically proposed Sotopia-Eval dimensions. We note that the other five dimensions in Sotopia are still valuable evaluation dimensions for social intelligence. However, behaviors captured by dimensions Financial and Material Benefits, Knowledge, and Secret are often also captured by Goal Completion and Believability. As a result, AutoLibra produces the single Goal Achievement and Outcome Effectiveness by minimizing redundancy. Whereas, Relationship and Social Rules captures long-tailed behaviors not captured by AutoLibra.
@@
Similarly, for web navigation tasks, AutoLibra also discovers metrics such as Access Barrier Handling, Error Recovery and Adjustment, Step Efficiency and Action Redundancy, and Navigation Accuracy, which much more closely reflect concrete agent behavior than the failure analysis categories proposed in previous work (He et al., 2024; Zhou et al., 2024c), where they are often simply classified as “navigation stuck”. We also find additional metrics that are not mentioned in the failure analysis, such as Query and Search Strategy Efficiency and Final Output and Summarization Quality, which are frequent issues (with frequencies of 7% and 18%). Since AutoLibra only observes the behavior of the agents, it cannot interpret the neural representation, not able to capture the visual grounding issues, which are mentioned in the WebVoyager paper. This further demonstrates AutoLibra’s utility in extracting behavior-salient metrics, and particularly demonstrates its ability to obtain fine-grained metrics that expert design would not have been able to extract.
@@ 5 AutoLibra as a ladder [Uncaptioned image]: Agent improvement with AutoLibra
Baba-Is-AI (5.1) WebVoyager (5.2)
Avg. Metrics Success Rate Avg. Metrics Success Rate
Iter 0 15.8 33.3 45.9 34.8
Iter 1 28.2 40.0 53.8 35.5
Iter 2 40.7 44.4 61.3 38.1
Iter 3 56.7 52.7 69.9 39.7
Table 3: Using AutoLibra-induced metrics as optimization targets for prompt engineering and finetuning leads to improvements in the average metric positive rates (Avg. Metrics) and task success rate (Success Rate) defined by the two environments respectively.
As AutoLibra can automatically induce metrics from human feedback, a natural question to ask is whether it can enable self-regulated improvement in agents through iterative feedback. In this section, we will study can AutoLibra be used a ‘ladder’, inducing metrics as new optimization targets in each iteration to improve agent performance. We investigate two methods for optimizing agents: prompt engineering and fine-tuning, across two domains: text game and web browsing. Across all tested methods and domains, AutoLibra induces new metrics in each iteration to cover emergent agent behaviors, and enables continuous improvements on both induced metrics and task success. Fig. 8 illustrates the results.
Figure 8: AutoLibra iteratively discovers new optimization targets throughout the agent optimization process. Each emoji pair represents an induced metric; the arrows and text between iterations represents the changes made to prompts informed by metrics. Purple boxes indicate the newly emerging metrics, and green arrows indicate significant improvements over the previous iteration. The metric names can be found in Tab. 4 and App. Tab. 10.
5.1 Can AutoLibra induce good optimization targets for prompt engineering?
To answer this question, we use Baba-Is-AI (Cloos et al., 2024; Paglieri et al., 2024) as a case study, since it is not yet saturated by state-of-the-art LLM agents. This game requires metacognitive skills, such as breaking and building new rules and planning ordered subtasks; a full list of game rules is listed in App. C. We use the top model on the leaderboard [^7] https://balrogai.com (which when we started the experiments, was GPT-4o) (OpenAI et al., 2024) for the agent policy, while we observe similar improvements in Claude-3.5-Sonnet [^8] https://www.anthropic.com/news/claude-3-5-sonnet as well.
To avoid overfitting to the tasks used in optimization, we fine-tune on 6 of 40 tasks in Baba-Is-AI (Paglieri et al., 2024), but report performance on all 40 tasks. In one Iteration, the human prompt engineers (two of the authors) generate 18 agent trajectories (3 trajectories / task), use AutoLibra to induce metrics, then optimize the prompt to achieve higher performance on the induced metrics without considering the task success rate, and start a new iteration when one or more metrics improved significantly (over 30%). The procedure and experiment setup are detailed in App. Alg. A and App. B.
Results
As shown in Fig. 8 and Tab. 3, the performance on Babaisai task increased from 25% to 55% between Iteration 0 and Iteration 3, with each iteration resulting in greater performance than the previous; this is replicated across the held-out tasks alone, (complete task results tabulated in App. D). This represents a substantial increase compared to the highest base model performance of 33% on Baba-Is-AI (Paglieri et al., 2024). With only 6 tasks used in inducing the metrics, the improvement on all 40 tasks showing the metrics induced by AutoLibra are generalizable to unseen tasks. Among induced metrics (App. Tab. 4), agent performance increases correspondingly to prompt changes targeting those metrics, an example being Win Condition Recognition [Uncaptioned image], which increases from 28% to 50% from Iterations 0-1 after few-shot prompting (prompt in App. G) is introduced to teach the identification of a win condition rule. The complete list of observations and reasoning for prompt engineering is in the App. H. With AutoLibra, prompt engineers could receive feedback on whether their prompt changes lead to desired behavior changes in agents.
Emoji It. Metric
[Uncaptioned image] 0 Win Condition Recognition
[Uncaptioned image] 0 Rule Modification
[Uncaptioned image] 0 Direct Navigation Efficiency
[Uncaptioned image] 0 Context-Sensitive Decision Making
[Uncaptioned image] 1 Win Rule Construction
[Uncaptioned image] 1 Selective Interaction w/ Relevant Obj.
[Uncaptioned image] 1 Rule Manipulation and Execution
[Uncaptioned image] 2 Subtask Coordination
[Uncaptioned image] 3 Immovable Interaction
Table 4: Metrics and Turn of Induction
for Baba-Is-AI
Qualitative observation of agent trajectories reveals behaviors commensurate with induced metric scores; the agent random-walks in Iteration 0, navigates towards specific objectives but gets stuck or trapped in a loop on long-horizon tasks in Iteration 1 and Iteration 2, and fully understands basic subtasks (atomic goals on the critical path to environment completion) and the correct order of subtasks to successfully complete an environment in Iteration 3. A per-metric score breakdown is listed in Appendix E, and a full per-iteration documentation of code changes and results is presented in Appendix H.
Generalization to other models When replacing GPT-4o with Claude-3.5 as the LLM used in the agent (Appendix D), performance was found to be similar across all iterations, with held-out task performance increasing by 20% to 55% between Iteration 0 and Iteration 3 and qualitatively similar trajectory performance and agent behaviors observed between the two LLMs. This demonstrates that the improvements realized by AutoLibra are generalizable to other LLMs, and that the induced metrics are robust to changes in the underlying LLM.
Generalization to other tasks To evaluate whether this pipeline can be generalized to other domains, we conduct the same pipeline to MiniHack (Samvelyan et al., 2021), an environment whose tasks and action space are more complex than Baba-Is-AI. Similarly to Baba-Is-AI, in three iterations, the task completion rate was observed to increase to 25%, an improvement of 15% versus the baseline of 10%, validating AutoLibra’s general utility in improving agent performance; full results are available in the Appendices I-N.
5.2 Can AutoLibra induce good optimization targets for agent fine-tuning?
To answer this question, we use web browsing (via the WebVoyager (He et al., 2024) benchmark) to compare the effectiveness of AutoLibra with existing goal-oriented LLM-as-a-Judge methods in rejection behavior cloning methods for agent fine-tuning. We use a diverse dataset (NNetNav-live) composed of agent trajectories of open-end exploration on 20 live websites (Murty et al., 2025). NNetNav-live tasks are assigned a task label post-hoc, to ensure they have no overlap with and are more complex than the WebVoyager evaluation datasets used in this subsection. Using the state-of-the-art agent Llama-8b-nnetnav-live, which was finetuned with trajectories selected with LLM-as-a-Judge instructed to evaluate trajectories based on goal completion, and following a similar pipeline to §5.1, we improve the agents for 3 iterations, obtaining feedback on 18 trajectories in each iteration. Instead of prompt engineering, we use the induced metrics to choose the sampled trajectories with top 10% sum scores of all metrics for fine-tuning the agent policy. Throughout the training procedure, we use the tasks in NNetNav-live (Murty et al., 2025) for metric induction and agent finetuning. This process yields a 4.9% task completion rate improvement on WebVoyager, shown in Fig. 8 and Tab. 3, while using naive LLM-as-a-Judge instructed to evaluate goal completion does not yield significant improvement in metrics or success rate. This shows that AutoLibra-induced metrics serve as better optimization targets than the goal-oriented evaluation metric alone; theseinduced metrics are shown in App. Tab. 10.
@@ 6 Related Work
Evaluating AI agents
Much of the work in AI agent evaluation focuses around benchmarks which contains both task suites and evaluation metrics. In addition to the datasets we used in this paper, SWE-Bench (Jimenez et al., 2024) uses human-written unit tests as evaluation metrics; Embodied Agent Interface (Li et al., 2024) provides fine-grained evaluation for LLM-based embodied agents; $\tau$-Bench (Yao et al., 2024) compares database states for evaluation; concurrent work AgentRewardBench (Lù et al., 2025) builds a benchmark for reward models for web agents. Recently, there are observatory tools including Galileo (Galileo, 2025), Vertex AI Gen AI (Cloud, 2025), and Docent (Meng et al., 2025) which provide user interfaces to visualize agent failure modes. Generating intrinsic rewards have also been studied in the reinforcement learning community (Du et al., 2019; Pathak et al., 2017; Laskin et al., 2022) to encourage exploration, sub-task completion, or skill discovery. In contrast to these, AutoLibra is a pure data-driven task-agnostic method without predefined failure taxonomy for generating interpretable metrics for agents.
Learning from natural language and human feedback
Researchers have been studying reinforcement learning with language feedback to provide a dense reward to agents (Goyal et al., 2019). Since LLM agents are even harder to train with sparse reward, there is substantial interest in training LLM agents from natural language feedback. Chen et al. (2024) propose an imitation learning method for learning from human feedback; Text2Reward (Xie et al., 2024) uses code generation to generate robot reward functions from open-ended human feedback; our work (Chen et al., 2025a) uses feedback to the improvement agent policy with prompting and then align the unprompted agent policy with the prompted one; Shi et al. (2024) propose a new model architecture to incorporate human feedback into policy learning. On the other hand, human non-open-ended feedback is also incorporated in training agents, including rating feedback (Nguyen et al., 2017), preference feedback (Christiano et al., 2017), demonstrative feedback (Shaikh et al., 2025). Unlike these papers, AutoLibra induces metrics from feedback from all annotated instances and generates metrics that are generalizable to different tasks and useful for both evaluation and agent fine-tuning.
Automatic thematic analysis
Thematic analysis is a powerful tool for qualitative study through coding and iterative creation of themes. Gauthier & Wallace (2022) provide computational tools to aid this process; Hong et al. (2022) and Gebreegziabher et al. (2023) explore human-AI collaboration in thematic analysis; LLooM (Lam et al., 2024), is an automatic method for concept induction, is closest to and influences our approach. This paper completes the loop of concept induction by using the meta-evaluation step to optimize the induced metrics, and apply this social science technique to agent evaluation.
@@ 7 Conclusion and Future Work
In this paper, we propose AutoLibra, a new paradigm for agent evaluation, one of the first works to explore adaptable trajectory-derived evaluation heuristics, offering substantial advantages in agent training over traditional end-to-end evaluation. We find that this framework is generalizable to a diverse range of agent tasks, provides new insights into agent behaviors, and identifies strong optimization targets for agent improvement. There are a few directions for further extending and applying this framework.
Behavior-centric evaluation AutoLibra leads a paradigm shift from end-to-end agent evaluation (analogous to “integration tests” in software development) to evaluation with granular metrics that measure agents’ concrete behaviors (analogous to “unit tests”). Future work can study whether this process can be improved through better human-AI collaboration.
Sub-trajectory feedback from humans In AutoLibra, we label each trajectory with one piece of feedback, and ground it into the agents’ concrete behavior which is at the sub-trajectory level. In the future, researchers can let users directly give feedback for one or multiple steps in the trajectory, which should lead to better feedback grounding results. Similarly, user feedback can be collected during the interaction instead of after the agent has completed the tasks, which is a more user-friendly way to gather high quality feedback data.
Wider exploration of agent improvement methods In this paper, we only explored two methods for agent improvement to show the utility of AutoLibra. In the future, researchers can use AutoLibra to provide dense rewards for individual steps, and use reinforcement learning to train LLM agents with these dense rewards.
Limitation and Ethics Statement
The scope of this paper is text-based agents, which does not include agents with multimodal observation or action spaces. Within human-aided experiments, we are also limited by the diversity of human annotators. The annotation of the data in this paper, except for the feedback collected by Shao et al. (2024), is done by the authors, who are experts in AI agents. We do not address the effect of the experience of the annotators on the induced metrics. Finally, induced metrics should be used with caution, these could reflect the internal biases of the LLMs used to extract them.
Acknowledgment

### v3

@@ 3.3 How aligned are the steps in AutoLibra with human judgment?
Grounding 0.95 0.95 0.98 0.93 0.93 0.95 ($\pm 0.03$)
LLM-as-a-Judge 0.90 0.85 0.95 1.00 0.90 0.92 ($\pm 0.04$)
Meta-Evaluation 0.98 0.90 0.85 0.83 0.95 0.90 ($\pm 0.04$)
@@ 4 AutoLibra as A Lens [Uncaptioned image] : Agent Evaluation with AutoLibra
CoGym
Table 2: AutoLibra-induced metrics and expert-proposed evaluation dimensions and failure categories. (Percentage %) denotes failure frequency or score from AutoLibra or the original papers.
@@
Sotopia (Zhou et al., 2024b) proposed 7 dimensions for evaluating social intelligence in AI agents. With AutoLibra, we recover the exact dimension Goal Completion, and 3 metrics as the subdimensions of Believability, indicating that Believability could be too high-level, while AutoLibra provides more concrete breakdown metrics. The failure rate (frequency of a score of -1 metric rating, indicating the agent performs poorly on that metric) measured by AutoLibra in these two categories roughly matches the score of the Sotopia dimensions of the agent we studied. AutoLibra induces another four metrics overlooked in the heuristically proposed Sotopia-Eval dimensions. We note that the other five dimensions in Sotopia are still valuable evaluation dimensions for social intelligence. However, behaviors captured by dimensions Financial and Material Benefits, Knowledge, and Secret are often also captured by Goal Completion and Believability. As a result, AutoLibra produces the single Goal Achievement and Outcome Effectiveness by minimizing redundancy. Whereas, Relationship and Social Rules captures long-tailed behaviors not captured by AutoLibra.
@@
Similarly, for web navigation tasks, AutoLibra also discovers metrics such as Access Barrier Handling, Error Recovery and Adjustment, Step Efficiency and Action Redundancy, and Navigation Accuracy, which much more closely reflect concrete agent behavior than the failure analysis categories proposed in previous work (He et al., 2024; Zhou et al., 2024c), where they are often simply classified as “navigation stuck”. We also find additional metrics that are not mentioned in the failure analysis, such as Query and Search Strategy Efficiency and Final Output and Summarization Quality, which are frequent issues (with frequencies of 7% and 18%). Since AutoLibra only observes the behavior of the agents, it cannot interpret the neural representation, not able to capture the visual grounding issues, which are mentioned in the WebVoyager paper. This further demonstrates AutoLibra’s utility in extracting behavior-salient metrics, and particularly its ability to obtain fine-grained metrics that expert design would not have been able to extract.
@@ 5 AutoLibra as A Ladder [Uncaptioned image] : Agent Improvement with AutoLibra
Figure 4: AutoLibra iteratively induce metrics and improves the agent prompts through optimizing for the induced metrics. Although not optimized for, the success rate of the agent continuously improve until Stage 3, when the agent begins to overthink.
As AutoLibra can automatically induce metrics from human feedback, a natural question to ask is whether it can enable self-regulated improvement in agents through iterative feedback. This can be achieved through optimizing the agent prompts towards higher scores on the metrics extracted by AutoLibra. To answer this question, we use a challenging 2D game Baba-Is-AI (Cloos et al., 2024; Paglieri et al., 2024) as a benchmark. Inspired by Baba-Is-You, this game requires not only following rules to achieve goals, but also manipulating the rules, even self-referential ones. For example, in the game illustrated in Fig. 5, the agent needs to change self-referential rules from baba is you, to door is you to control the green door on the other side of the wall, form a new win rule ball is win, and navigate to the red ball to achieve the win condition. To achieve a high score on this dataset, the agent needs not only planning, but also metacognitive skills, which is very challenging for LLM agents with frontier models as shown in the Balrog benchmark (Paglieri et al., 2024). In this experiment, we use Gemini-2.5-Flash (Team et al., 2025) for the agent, AutoLibra, and agent prompt optimization, throughout the experiment, which will be referred as the LLM in this section. Gemini-2.5-Flash is ranked as the 3rd place, with a success rate of $50.8\%\pm 4.6\%$ on the Balrog leaderboard for Baba-is-AI at the time of submission, and the state-of-the-art result is $56.7\%\pm 4.5\%$. We chose this model due to the tradeoff between the cost and the performance.
Figure 5: Example of Baba-Is-AI game.
Fig. 4 illustrated our procedure, and summarized the results. In this experiment, we employ an iterative process by improving the agents in 3 stages through providing human feedback on 6 out of 40 tasks in the Baba-Is-AI (Paglieri et al., 2024) benchmark. In this way, we can study if the feedback provided for training tasks can be generalized to unseen tasks.
Before each stage we show human annotators 3 trajectories per task for the 6 tasks, gather the feedback, and apply AutoLibra iterative metric induction process (§3.2). This results in 5 metrics for Stage 1 and 2, and another 1 metric for Stage 3. Within each stage, we iteratively feed 1 LLM agent trajectory on each of these 6 tasks, together with evaluation results based on these AutoLibra-induced metrics to the LLM to improve the prompt of the LLM agent. This process results in continuous improvement not only on the running maximum metric scores, the cumulative average metrics, but also game success rate. Fig. 4 shows these statistics on the whole 40 tasks, although we only use 6 out of the 40 tasks in the whole optimization process. Upon examining the agent trajectories, we find the skills learned in the process. In the first stage, the agent learns to find rules to form based on the map boundary, which could be a result of an induced metric map-n-constraint-recognition. Similarly, more advanced skills are learned in Stage 2 and 3, including forming win conditions and self-referential rules, probably as a result of metric rule-manipulation-proficiency.
Our results show that the metrics induced by AutoLibra form effective objectives for improving the agents through prompt optimization. It should note that AutoLibra is a metric induction method, which is orthogonal to learning algorithms, including prompt optimization, fine-tuning or reinforcement learning. We show that this process improves agent success rate by 20% without optimizing for success rate, and in the future, researchers can study the effect of employing other learning algorithm.
@@ 6 Related Work
Evaluating AI agents Much of the work in AI agent evaluation focuses around benchmarks which contains both task suites and evaluation metrics. In addition to the datasets we used in this paper, SWE-Bench (Jimenez et al., 2024) uses human-written unit tests as evaluation metrics; Embodied Agent Interface (Li et al., 2024) provides fine-grained evaluation for LLM-based embodied agents; $\tau$-Bench (Yao et al., 2024) compares database states for evaluation; concurrent work AgentRewardBench (Lù et al., 2025) builds a benchmark for reward models for web agents. Recently, there are observatory tools including Galileo (Galileo, 2025), Vertex AI Gen AI (Cloud, 2025), and Docent (Meng et al., 2025) which provide user interfaces to visualize agent failure modes. Generating intrinsic rewards have also been studied in the reinforcement learning community (Du et al., 2019; Pathak et al., 2017; Laskin et al., 2022) to encourage exploration, sub-task completion, or skill discovery. In contrast to these, AutoLibra is a pure data-driven task-agnostic method without predefined failure taxonomy for generating interpretable metrics for agents.
Learning from natural language and human feedback Researchers have been studying reinforcement learning with language feedback to provide a dense reward to agents (Goyal et al., 2019). Since LLM agents are even harder to train with sparse reward, there is substantial interest in training LLM agents from natural language feedback. Chen et al. (2024) propose an imitation learning method for learning from human feedback; Text2Reward (Xie et al., 2024) uses code generation to generate robot reward functions from open-ended human feedback; our work (Chen et al., 2025) uses feedback to the improvement agent policy with prompting and then align the unprompted agent policy with the prompted one; Shi et al. (2024) propose a new model architecture to incorporate human feedback into policy learning. On the other hand, human non-open-ended feedback is also incorporated in training agents, including rating feedback (Nguyen et al., 2017), preference feedback (Christiano et al., 2017), demonstrative feedback (Shaikh et al., 2025). Unlike these papers, AutoLibra induces metrics from feedback from all annotated instances and generates metrics that are generalizable to different tasks and useful for both evaluation and agent fine-tuning.
Thematic analysis Thematic analysis is a powerful tool for qualitative study through coding and iterative creation of themes. Gauthier & Wallace (2022) provide computational tools to aid this process; Hong et al. (2022) and Gebreegziabher et al. (2023) explore human-AI collaboration in thematic analysis; LLooM (Lam et al., 2024), an automatic method for concept induction, closly aligns with and informs our approach. This paper completes the loop of concept induction by using the meta-evaluation step to optimize the induced metrics, and apply it to agent evaluation.
@@ 7 Conclusion and Future Work
This work introduces AutoLibra, a new paradigm for agent evaluation, one of the first works to explore adaptable trajectory-derived evaluation heuristics, offering substantial advantages in agent training over traditional end-to-end evaluation. We find that this framework is generalizable to a diverse range of agent tasks, provides new insights into agent behaviors, and identifies strong optimization targets for agent improvement. There are a few directions for further extending and applying this framework. (1) Behavior-centric evaluation AutoLibra leads a paradigm shift from end-to-end agent evaluation (analogous to “integration tests” in software development) to evaluation with granular metrics that measure agents’ concrete behaviors (analogous to “unit tests”). Future work can study whether this process can be improved through better human-AI collaboration. (2) Sub-trajectory feedback from humans In AutoLibra, we label each trajectory with one piece of feedback, and ground it into the agents’ concrete behavior which is at the sub-trajectory level. In the future, researchers can let users directly give feedback for one or multiple steps in the trajectory, which should lead to better feedback grounding results. Similarly, user feedback can be collected during the interaction instead of after the agent has completed the tasks, which is a more user-friendly way to gather high quality feedback data. (3) Wider exploration of agent improvement methods In this paper, we only explored non-parametric for agent improvement to show the utility of AutoLibra. Future work can use AutoLibra to provide dense rewards for individual steps, and use reinforcement learning to train agents with these dense rewards.
8 The Bigger Picture
Humans learn not only from their own experience, but crucially from collective knowledge transmitted through social interaction (Tomasello et al., 1993). This capacity for social learning and teaching enables knowledge accumulation across generations, allowing us to build upon rather than reinvent skills and artifacts (Humphrey, 1976). At the core of this capability lies a distinctive form of intelligence that sets humans apart from species lacking cumulative culture: social intelligence. A fundamental question emerges: how can we leverage this social intelligence to enable AI agents to learn from humans and potentially teach them in similar ways?
Current training paradigms for language agents predominantly focus on behavior cloning or reinforcement learning from self-generated experience. However, behavior cloning has inherent limitations for skill acquisition. First, it requires teachers or demonstrators to operate in the same environment as the learning agent, constraining the scalability of knowledge transfer. Second, agents learning from one or a few demonstrations may develop misunderstandings about what constitutes appropriate behavior. In such cases, explicit verbal feedback from teachers becomes crucial for establishing the boundaries of acceptable behaviors and clarifying underlying principles that cannot be easily inferred from demonstrations alone.
AutoLibra offers a computational approach to address these limitations by incorporating the verbal feedback that humans naturally exchange in daily interactions to systematically improve agent behavior. Looking forward, the broader vision is to deploy AI agents in real-world settings where they can learn from diverse human feedback at scale. This approach could enable agents to acquire skills from the collective experience of all users, forming both general competencies and specialized skill sets at personal and organizational levels, mirroring how human communities accumulate and share knowledge.
Going one step further, such agents could serve as intermediaries for accelerating skill and culture transmission between humans themselves—distilling expertise from proficient users into interpretable metrics and actionable feedback that can guide novices, thereby compressing traditional learning curves and democratizing access to specialized knowledge.
Acknowledgments